Find the area of the region bounded by the parabola  y=xand the line x=4y-2 ; explain briefly .

Grade:

1 Answers

Sarbajit Mazumdar
5 years ago
first solve two equations to get $x=\frac{1}{2},-\frac{1}{4}$ these are mainly cot points of two curves 
now simply find the area as $A=\int{-\frac{1}{4}}{\frac{1}{2}}(\frac{x+2}{4}-x^{2})dx$ to get the answer.